OpenStudy (anonymous):

Use the zeros of the following quadratic to find the x-value of the vertex. y = x^2 – 4x – 5 A. 5 B. 2 C. –2 D. –5

OpenStudy (anonymous):

Lol , Im going to help you even though we are sitting next to each other !

OpenStudy (anonymous):

Vertex = 2

OpenStudy (anonymous):

when we are solving the quadratic equations then we have to take it equals to zero. that is we have to take 0 = x^2 – 4x – 5 now we have to solve this for x. we use here AC method of factoring. The AC method of factoring is used to factor polynomials of the form Ax² + Bx + C. A, B, and C represent constants and x is the variable. Determine AC by multiplying the A term and C term. x^2 – 4x – 5 by comparing we get A = 1 B = - 4 and C = -5 so the answer is A . Find two numbers that add to the B term and multiply to AC. that is find two numbers that add to -4 and multiply to AC that is -5*1=-5 so here we get -5+1 = -4 and when we multiply them we get -5*1 = -5 Call the smaller number M and the larger number N. so the smaller number is -5 and larger number = 1 Rewrite the original equation as Ax² + Mx + Nx + C. that is x^2 -5x +x -5 now factor out the common term from first two terms. here x is the common term in first two terms so that we get x(x - 5) + x - 5 also find the common factor from last two terms here 1 is common factor from last two terms. so that we get x (x-5) +1(x - 5) now we apply the distributive property. (x+1)(x-5) we get this factors. in the original problm we have 0 = x^2-4x-5 we get right side is (x+1)(x-5) so that 0 = (x+1)(x-5) so we write it like (x+1) = 0 or (x-5) = 0 from first factor we get x = -1 and x = 5